OpenStudy (nykoreyan):

Solve the system of equations using substitution. y = x + 6 y = –2x – 3 a.(4, –11) b.(-6,3/2 c.(–3, 3) d.(1, 7)

OpenStudy (nykoreyan):

@aloud @JasonSpade @ganeshie8 @imqwerty

OpenStudy (anonymous):

Solve y=x;6y=−2x−3 Steps: I will try to solve your system of equations. y=x;6y=−2x−3 Step: Solvey=xfor y: Step: Substitutexforyin6y=−2x−3: 6y=−2x−3 6x=−2x−3 6x+2x=−2x−3+2x(Add 2x to both sides) 8x=−3 8x8=−38(Divide both sides by 8) x=−38 Step: Substitute−38forxiny=x: y=x y=−38 Answer: y=−3/8 and x=−3/8

OpenStudy (anonymous):

C. (-3,3)

OpenStudy (nykoreyan):

^^^^ thank you man ........i dont ask for explanations ..i ask for a b c or d thats what gets u a medal and fan